OpenStudy (anonymous):

What are the sine, cosine, and tangent of 5 pi over 4 radians? i already know how to convert radians to degrees but im not sure where to go from there

OpenStudy (anonymous):

so you have 225 degrees?

OpenStudy (anonymous):

and do you know how to use the unit circle?

OpenStudy (anonymous):

Kind of. I couldn't quite grasp the concept of it

OpenStudy (anonymous):

ok, the unit circle is just a circle graph, it shows you the exact values of certain points at certain degress on the way around

OpenStudy (anonymous):

so like with 225 degrees its going to read as \[ \frac{ 5\pi }{ 4 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So then how do i take that and put it in Sin, Cosine, and Tangent?

OpenStudy (anonymous):

idk if you are this far, but sin=y, cos=x, and tan=y/x

OpenStudy (anonymous):

so you are given the coordinates for 225 on the graph, and you just use them as an ordered pair

OpenStudy (anonymous):

idk what just happened... ignore that please...

OpenStudy (anonymous):

mmmkay, hope it helped

OpenStudy (anonymous):

wait so if i have the coordinate points then i have two x's and two y's so which ones do i use and what do i do with them?

OpenStudy (anonymous):

why would do you have two x's an y's?

OpenStudy (anonymous):

pssshhhh my bad. im blind so then sine and cosine are both \[\frac{ -\sqrt{2} }{ 2 }\] and\[\frac{ -\sqrt{2} }{ 2 }*\frac{ 2 }{ -\sqrt{2} }\] is tangent

OpenStudy (anonymous):

i got it :) thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

and good job figuring it out